# On estimates of trace-norm distance between quantum Gaussian states

### v5

@@ Abstract
In the paper [1] estimates for the trace-norm distance between two quantum Gaussian states of CCR in terms of the mean vectors and covariance matrices were derived and used to evaluate the number of elements in the $\varepsilon-$net in the set of energy-constrained Gaussian states. In the present paper we obtain different estimates based on a fidelity-like quantity which we call states overlap [2]. Our proof is more direct leading to estimates which are sometimes even more stringent, especially in the cases of pure or gauge-invariant states. They do not depend on number of modes and hence can be extended to the case of bosonic field with infinite number of modes. These derivations are not aimed to replace the useful inequalities from [1]; they just show an alternative approach to the problem leading to different results. In the Appendix we briefly recall our results from [3] concerning estimates of the overlap for general fermionic Gaussian states of CAR.
@@ 1 Introduction and results
A hypothetical proof based on the fidelity $\mathrm{Tr}\left|\sqrt{\rho_{1}}\sqrt{\rho_{2}}\right|$ indeed looks to be too difficult. In the present paper, instead of the fidelity we try a simpler quantity $\mathrm{Tr}\sqrt{\rho_{1}}\sqrt{\rho_{2}}$, which we call quantum states overlap, and use the corresponding estimate of the trace-norm distance from our paper [2]. This allows us for a more straightforward proof of the estimates which are sometimes even more stringent, especially in the cases of pure or gauge-invariant states. They do not depend on number of modes and hence can be extended to the case of bosonic field with infinite number of modes. However these derivations are not aimed to replace the useful inequalities from [1] which are adjusted for the purposes of that important paper; they just show a possibility of an alternative approach to the problem leading to different estimates.
@@
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq\sqrt{8\min\left(a,b\right)\left\|m\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (1)
@@
$\left\|\rho_{\alpha}-\rho_{{}_{\beta}}\right\|_{1}\leq 4\sqrt{\left(a+b\right)\left\|\alpha-\beta\right\|_{1}-\left\|\alpha-\beta\right\|_{2}^{2}}.$ (2)
@@
$2\sqrt{4\min\left(a,b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}+\min\left\{P(a,b),P(b,a)\right\}\left\|\alpha-\beta\right\|_{1}},$
@@
The proofs are presented in sec. 3. In the concluding sec. 4 we compare our estimates with those obtained in [1]. In the Appendix we briefly recall our results from [3] concerning estimates of overlap for general fermionic Gaussian states of CAR.
@@ A Gaussian state $\rho_{m,\alpha}$ is determined by its quantum characteristic function
Our approach is based on the second inequality in
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{8}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)}.$ (33)
@@
$\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)$
$=$ $4\mathrm{Tr}[\left(\alpha-\beta\right)+\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)]\Delta^{-1}[\left(\alpha-\beta\right)-\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)]\Delta^{-1}$
$=$ $4\mathrm{Tr}[\left(\alpha-\beta\right)\Delta^{-1}]^{2}-4\mathrm{Tr}[\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)\Delta^{-1}]^{2}$ (35)
$\leq$ $4\mathrm{Tr}[\alpha-\beta]^{2}+4\mathrm{Tr}[\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}]^{2},$ (36)
@@ 3.1 Pure states
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{\frac{1}{2}m^{t}\left(\alpha+\beta\right)^{-1}m+\frac{1}{4}\mathrm{Tr}\left(\alpha-\beta\right)^{2}}.$
Since $\left(\alpha+\beta\right)^{-1}\leq\alpha^{-1},\left(\alpha+\beta\right)^{-1}\leq\beta^{-1},$ we have by using (25)
$m^{t}\left(\alpha+\beta\right)^{-1}m$$\leq$ $\min\left(\left\|\alpha^{-1}\right\|,\left\|\beta^{-1}\right\|\right)\left\|m\right\|_{2}^{2}$ (38)
$\leq$ $4\min\left(a,b\right)\left\|m\right\|_{2}^{2},$
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq\sqrt{8\min\left(a,b\right)\left\|m\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (39)
@@ 3.2 Gauge-invariant states
$-4\mathrm{Tr}[\alpha-\beta]^{2}+4\mathrm{Tr}[\sqrt{\alpha^{2}-1/4}-\sqrt{\beta^{2}-1/4}]^{2}.$ (41)
@@
$4\left\|\alpha^{2}-\beta^{2}\right\|_{1}$ $=$ $4\left\|\left(\alpha-\beta\right)\circ\left(\alpha+\beta\right)\right\|_{1}$
$\leq$ $4\left(\left\|\alpha\right\|+\left\|\beta\right\|\right)\left\|\alpha-\beta\right\|_{1}$
$=$ $4\left(a+b\right)\left\|\alpha-\beta\right\|_{1},$
where $\circ$ denoted the Jordan product. Inserting in (41) we obtain the final estimate (2) for the trace-norm distance of gauge-invariant states.
@@ 5 Appendix: the case of fermionic Gaussian states of CAR
Here we briefly recall some results from our paper [3] where the overlap was computed for general fermionic Gaussian states of canonical anticommutation relations (CAR).
@@
Lemma 1. Let $A,B$ be operators in $H$ satisfying (50), then the overlap of $\rho_{A},\rho_{B}$ is
@@
Introducing $s(A,B)=\mathrm{Tr}\left(I-S^{\ast}S\right)$ we have the following estimates for the overlap:

### v6

@@ Abstract
In the paper [1] estimates for the trace-norm distance between two quantum Gaussian states of CCR in terms of the mean vectors and covariance matrices were derived and used to evaluate the number of elements in the $\varepsilon-$net in the set of energy-constrained Gaussian states. In the present paper we obtain different estimates based on a fidelity-like quantity which we call states overlap [2]. Our proof is more direct leading to estimates which are sometimes even more stringent, especially in the cases of pure or gauge-invariant states. They do not depend on number of modes and hence can be extended to the case of bosonic field with infinite number of modes. These derivations are not aimed to replace the useful inequalities from [1]; they just show an alternative approach to the problem leading to different results. In the Appendix we briefly recall our results from [3] concerning estimates of the states overlap for general fermionic Gaussian states of CAR.
@@ 1 Introduction and results
A hypothetical proof based on the fidelity $\mathrm{Tr}\left|\sqrt{\rho_{1}}\sqrt{\rho_{2}}\right|$ indeed looks to be too difficult. In the present paper, instead of the fidelity we try a simpler quantity $\mathrm{Tr}\sqrt{\rho_{1}}\sqrt{\rho_{2}}$, which we call here quantum states overlap [^1] Some authors call the overlap a different quantity $\mathrm{Tr}\rho_{1}\rho_{2}$., and use the corresponding estimate of the trace-norm distance from our paper [2]. This allows us for a more straightforward proof of the estimates which are sometimes even more stringent, especially in the cases of pure states. They do not depend on number of modes and hence can be extended to the case of bosonic field with infinite number of modes. However these derivations are not aimed to replace the useful inequalities from [1] which are adjusted for the purposes of that important paper; they just show a possibility of an alternative approach to the problem leading to different estimates [^2] After the first posting of the present paper, the sequel [9] of [1] appeared where the estimates of [1] were substantially improved by using still a different gradient method. Alternative estimates based on quantum Pinsker inequality targeted to Hamiltonian learning were demonstrated in [10]..
@@
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq\sqrt{2\left(a+b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (1)
@@
$\left\|\rho_{\alpha}-\rho_{{}_{\beta}}\right\|_{1}\leq\sqrt{2[\left(a+b\right)\left\|\alpha-\beta\right\|_{1}-\left\|\alpha-\beta\right\|_{2}^{2}]}.$ (2)
@@
$2\sqrt{\left(a+b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}+\min\left\{P(a,b),P(b,a)\right\}\left\|\alpha-\beta\right\|_{1}},$
@@
The proofs are presented in sec. 3. In the concluding sec. 4 we compare our estimates with those obtained in [1]. In the Appendix we briefly recall our results from [3] concerning estimates of state overlap for general fermionic Gaussian states of CAR.
@@ A Gaussian state $\rho_{m,\alpha}$ is determined by its quantum characteristic function
Our approach uses the second inequality in
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{2}\mathrm{Tr\,}d(\alpha,\beta)}.$ (33)
@@
$\mathrm{Tr\,}d(\alpha,\beta)=\frac{1}{4}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)$
$=$ $\mathrm{Tr}[\left(\alpha-\beta\right)+\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)]\Delta^{-1}[\left(\alpha-\beta\right)-\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)]\Delta^{-1}$
$=$ $\mathrm{Tr}[\left(\alpha-\beta\right)\Delta^{-1}]^{2}-\mathrm{Tr}[\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)\Delta^{-1}]^{2}$ (35)
$\leq$ $\mathrm{Tr}[\alpha-\beta]^{2}+\mathrm{Tr}[\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}]^{2},$ (36)
@@ 3.1 Pure states
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq\sqrt{m^{t}\left(\frac{\alpha+\beta}{2}\right)^{-1}m+\mathrm{Tr}\left(\alpha-\beta\right)^{2}}.$
Since $\sigma=\frac{\alpha+\beta}{2}$ satisfies (8) along with $\alpha,\beta$, it satisfies also (24), hence
$m^{t}\sigma^{-1}m\leq\left\|\sigma^{-1}\right\|\left\|m\right\|_{2}^{2}\leq 4\|\sigma\|\left\|m\right\|_{2}^{2}\leq 2(a+b)\left\|m\right\|_{2}^{2},$ (38)
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq\sqrt{2\left(a+b\right)\left\|m\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (39)
@@ 3.2 Gauge-invariant states
$\mathrm{Tr\,}d(\alpha,\beta)=-\mathrm{Tr}[\alpha-\beta]^{2}+\mathrm{Tr}[\sqrt{\alpha^{2}-1/4}-\sqrt{\beta^{2}-1/4}]^{2}.$ (41)
@@
$\left\|\alpha^{2}-\beta^{2}\right\|_{1}$ $=$ $\left\|\left(\alpha-\beta\right)\circ\left(\alpha+\beta\right)\right\|_{1}$
$\leq$ $\left(\left\|\alpha\right\|+\left\|\beta\right\|\right)\left\|\alpha-\beta\right\|_{1}$
$=$ $\left(a+b\right)\left\|\alpha-\beta\right\|_{1},$
where $\circ$ denoted the Jordan product. Inserting in (41) and using (33) we obtain the final estimate (2) for the trace-norm distance of gauge-invariant states.
@@ 5 Appendix: the case of fermionic Gaussian states of CAR
Here we briefly recall some results from our paper [3] where the states overlap was computed for general fermionic Gaussian states of canonical anticommutation relations (CAR).
@@
Lemma 1. Let $A,B$ be operators in $H$ satisfying (50), then the states overlap of $\rho_{A},\rho_{B}$ is
@@
Introducing $s(A,B)=\mathrm{Tr}\left(I-S^{\ast}S\right)$ we have the following estimates for the states overlap: